If the contribution margin ratio is 0.40, targeted operating income is $95,000, and targeted sales volume in dollars is $520,000, then the degree of operating leverage is ________.

3.28 times

0.46 times

1.50 times

2.19 times

Sales of Blistre Autos are 350,000, variable cost is 210,000, fixed cost is 90,000 tax rate is 40%. Calculate the operating leverage of the company.

1.50 times

2.80 times

4.67 times

1.80 times

Tony Manufacturing produces a single product that sells for $80. Variable costs per unit equal $45. The company expects total fixed costs to be $80,000 for the next month at the projected sales level of 2500 units. In an attempt to improve performance, management is considering a number of alternative actions. Each situation is to be evaluated separately.Suppose management believes that a $90,000 increase in the monthly advertising expense will result in a considerable increase in sales. Sales must increase by ________ to justify this additional expenditure?

2260 units

1125 units

2572 units

2286 units

Tony Manufacturing produces a single product that sells for $120. Variable costs per unit equal $50. The company expects total fixed costs to be $81,000 for the next month at the projected sales level of 3000 units. In an attempt to improve performance, management is considering a number of alternative actions. Each situation is to be evaluated separately.Suppose that management believes that a 11% reduction in the selling price will result in a 11% increase in sales. If this proposed reduction in selling price is implemented ________.

operating income will decrease by $39,600

operating income will increase by $18,744

operating income will increase by $20,856

operating income will decrease by $20,856
Answer #1

1. Contribution margin = 520000*0.40 = $208000

Degree of operating leverage = contribution margin /net operating income

= 208000/95000 = 2.19 times

2. Contribution margin = 350000-210000 = 140000

Net operating income = 140000-90000)*(1-0.40) = 30000

Operating leverage = 140000/30000 = 4.67 times